OpenStudy (anonymous):

how do you do √50x - √8x wrote the question wrong

OpenStudy (radar):

\[\sqrt{50x}-\sqrt{8x}\]\[\sqrt{25 * 2x}-\sqrt{4 * 2x}\]\[5\sqrt{2x}-2\sqrt{2x}\]\[3\sqrt{2x}\]

OpenStudy (anonymous):

thnx!